# Diversifying Robot Locomotion Behaviors with Extrinsic Behavioral Curiosity

### v3

@@ Abstract
Imitation learning (IL) has shown promise in robot locomotion but is often limited to learning a single expert policy, constraining behavior diversity and robustness in unpredictable real-world scenarios. To address this, we introduce Quality Diversity Inverse Reinforcement Learning (QD-IRL), a novel framework that integrates quality-diversity optimization with IRL methods, enabling agents to learn diverse behaviors from limited demonstrations. This work introduces Extrinsic Behavioral Curiosity (EBC), which allows agents to receive additional curiosity rewards from an external critic based on how novel the behaviors are with respect to a large behavioral archive. To validate the effectiveness of EBC in exploring diverse locomotion behaviors, we evaluate our method on multiple robot locomotion tasks. EBC improves the performance of QD-IRL instances with GAIL, VAIL, and DiffAIL across all included environments by up to 185%, 42%, and 150%, even surpassing expert performance by 20% in Humanoid. Furthermore, we demonstrate that EBC is applicable to Gradient-Arborescence-based Quality Diversity Reinforcement Learning (QD-RL) algorithms, where it substantially improves performance and provides a generic technique for diverse robot locomotion. The source code of this work is provided at https://github.com/vanzll/EBC.
@@ 1 Introduction
  • Our framework can potentially enhance any IRL application that requires learning diverse policies, and can also significantly improve existing SOTA QD-RL algorithm, thus providing a generic framework for future research of Diverse Robot Locomotion.
@@ 2.1 Quality Diversity Reinforcement Learning
Distinct from traditional policy optimization which aims to find a single policy to maximize the cumulative reward, Quality Diversity Reinforcement Learning (QD-RL) aims to find a set of high-performing and diverse policies. Parametrizing policies based on $\theta\in\Theta\subset\mathbb{R}^{n}$, the objective function (i.e., fitness) is defined as $f:\Theta\rightarrow\mathbb{R}$, which is the cumulative reward in RL context. The diversity is based on a $k$-dimensional measure function $m:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$. The goal is to find policies $\theta\in\Theta$ for each local region in the behavior space $B=\mathbf{m}(\Theta)$. In grid-based algorithms MAP-Elites, one discretizes $B$ into $M$ cells (local regions) and forms a policy archive $\mathcal{A}$ , where each cell $i=1,\dots,M$ represents a small hypercube $[\mathbf{a_{i}},\mathbf{b_{i}}]$ within a multi-dimensional grid of the behavior space. By filling new cells and replacing existing solutions if they are outperformed, the algorithm gradually evolves an archive of high-quality and diverse policies (Chatzilygeroudis et al., 2021; Pugh et al., 2016; Mouret & Clune, 2015a, b). Based on MAP-Elites, Co- variance Matrix Adaptation MAP-Elites (CMA-ME) (Fontaine et al., 2020) further adopts covariance matrix adaptation as the evolution algorithm to find new high-performing solutions more effectively and efficiently. In such algorithms, QD-RL can be treated equivalently to solving $M$ policy-optimization problems of the form
@@
[…]
@@ 3 Extrinsic Behavioral Curiosity for QD-IRL
3:for $\text{iter}\leftarrow 1$ to $N$ do
4:     $f,g,\nabla g,\mathbf{m},\nabla\mathbf{m}\leftarrow\text{VPPO.jacobian}(\theta_{\mu},\mathcal{R},\mathbf{m},N_{1})$triangleright approx grad using $\mathcal{R}$
@@
[…]
@@
8:         $c\sim\mathcal{N}(\mu,\Sigma)$triangleright sample gradient coefficients
@@
12:         $\Delta_{i}\leftarrow\text{update\_archive}(\theta_{i}^{\prime},f^{\prime},\mathbf{m}^{\prime})$triangleright get archive improvement of each solution.
@@
16:     $g^{\prime}(\theta_{\mu}):=c_{\mu,0}g(\theta_{\mu})+\sum_{j=1}^{k}c_{\mu,j}m_{j}(\theta_{\mu})$, where $c_{\mu}=\mu^{\prime}$triangleright construct EBC-QD objective function
17:     $\theta_{\mu}^{\prime}\leftarrow\text{VPPO.train}(\theta_{\mu},g^{\prime},N_{2},\mathcal{R})$triangleright walk search policy using reward model $\mathcal{R}$
18:     .$\mathcal{R}.\text{update}(\mathcal{D},\theta_{\mu}^{\prime})$triangleright update reward model
@@ 3.3.3 DiffAIL-EBC
$=\exp\left(-||\epsilon-\epsilon_{\psi}(\sqrt{\alpha_{i}}x_{t}+\sqrt{1-\alpha_{t}}\epsilon,t)||^{2}\right),$
@@ 4.1 Learning Diverse High-Performing Skills from Limited Demonstrations
Experiments on Various Measure Metrics To further assess the generality of our framework, we applied EBC to the GAIL baseline (GAIL-EBC) and evaluated it on tasks that differ in both the type and the number of dimension of the QD measure $\mathbf{m}$. Specifically, we used: 1) Jump ($dim{=}1$), a scalar that records the height of the lowest foot at each step; 2) Angle ($dim{=}2$), a 2-D vector $\phi_{t}=(\cos\alpha,\ \sin\alpha)^{\top}$ capturing the yaw orientation $\alpha$ of the torso about the $z$-axis; and 3) Feet Contact ($dim{=}4$), a 4-D vector that stores, for each of the ant’s four legs, the fraction of time steps during which the leg touches the ground. Table 1 reports that GAIL-EBC improves QD-Score of GAIL throughout every scenario and significantly improve the Coverage of GAIL in humanoid environment.
@@
Environment Measure ($dim$) Model QD-Score Coverage (%)
@@
GAIL-EBC 90 631$\pm$ 61 894 85.0$\pm$ 15.0
Ant Feet Contact (4) GAIL --484 468 $\pm$ 3 264 75.84 $\pm$ 0.0
GAIL-EBC --52 521$\pm$ 383 213 77.28$\pm$ 3.2
@@ 5 Conclusion
In this work, we propose Extrinsic Behavioral Curiosity (EBC), a simple but effective approach which is highly extendable, and theoretically derived the rationale behind it. EBC can be seamlessly integrated into and can potentially improve any IRL method in QD contexts, providing the a generic QD-IRL framework for future research. Our framework follows the paradigm of IRL to learn a QD-enhanced reward function, and uses a QD-RL algorithm to optimize a policy archive. By encouraging behavior-level exploration, our framework addresses the key challenges of QD-IRL. Extensive experiments show that our framework can often improve QD-IRL methods and significantly outperform baselines. Additionally, we empirically prove that EBC can also significantly improve the performance of the state-of-the-art QD-RL algorithm, PPGA, highlighting the sub-optimality of the true reward function in QD-RL settings. With EBC, Gradient-Arborescence based QD-RL and QD-IRL algorithms can learn archives of diverse and high-performing policies more efficiently, pushing the boundary of diverse robot locomotion. We also provide potential limitations of our work in Appendix K and additional discussion in Appendix L.
@@ E.2 Hyperparameters for IL
Variance schedule ($\beta_{t}$) linear from $41e-$ 4 to $22e-$ 2
@@ Appendix G QD Metrics Performance Results
TrueReward $6.54\times 10^{6}\pm 1.47\times 10^{5}$ 94.68$\pm$1.11 8,724$\pm$384 2,764$\pm$72 $3.66\times 10^{6}\pm 3.20\times 10^{4}$ 77.88$\pm$1.19 5,454$\pm$189 1,882$\pm$12 $7.38\times 10^{6}\pm 2.37\times 10^{6}$ 62.13$\pm$17.22 9,636$\pm$78 4,700$\pm$185
DiffAIL-EBC (Ours) $3.99\times 10^{6}\pm 3.11\times 10^{5}$ 100.0$\pm$0.0 4,290$\pm$117 1,598$\pm$125 $2.93\times 10^{6}\pm 7.59\times 10^{4}$ 88.63$\pm$1.06 2,283$\pm$109 1,320$\pm$23 $8.92\times 10^{6}\pm 6.60\times 10^{5}$ 93.29$\pm$0.41 5,598$\pm$298 3,824$\pm$268
DiffAIL $4.02\times 10^{6}\pm 5.82\times 10^{4}$ 99.72$\pm$0.28 3,646$\pm$41 1,614$\pm$28 $1.71\times 10^{6}\pm 4.08\times 10^{5}$ 52.14$\pm$0.74 1,989$\pm$646 1,304$\pm$294 $3.56\times 10^{6}\pm 3.71\times 10^{5}$ 76.32$\pm$0.48 5,151$\pm$132 1,869$\pm$206
VAIL-EBC (Ours) $3.78\times 10^{6}\pm 7.69\times 10^{4}$ 99.14$\pm$0.50 3,901$\pm$840 1,526$\pm$39 $3.19\times 10^{6}\pm 2.15\times 10^{5}$ 91.45$\pm$1.62 3,802$\pm$375 1,398$\pm$72 $7.26\times 10^{6}\pm 3.10\times 10^{5}$ 97.63$\pm$0.59 5,824$\pm$489 2,975$\pm$145
VAIL $3.62\times 10^{6}\pm 4.00\times 10^{4}$ 91.68$\pm$3.80 5,123$\pm$401 1,582$\pm$83 $2.40\times 10^{6}\pm 2.13\times 10^{5}$ 71.40$\pm$2.70 3,570$\pm$417 1,341$\pm$73 $5.09\times 10^{6}\pm 6.86\times 10^{5}$ 65.61$\pm$2.77 7,056$\pm$250 3,094$\pm$304
GAIL-EBC (Ours) $2.64\times 10^{6}\pm 9.21\times 10^{4}$ 98.13$\pm$1.69 2,770$\pm$440 1,074$\pm$25 $3.42\times 10^{6}\pm 1.36\times 10^{5}$ 91.49$\pm$0.98 4,022$\pm$569 1,493$\pm$48 $5.31\times 10^{6}\pm 5.78\times 10^{5}$ 98.00$\pm$0.23 5,140$\pm$852 2,166$\pm$241
GAIL $2.02\times 10^{6}\pm 8.36\times 10^{5}$ 67.83$\pm$16.05 5,115$\pm$218 1,165$\pm$341 $2.47\times 10^{6}\pm 2.88\times 10^{5}$ 69.27$\pm$4.51 4,031$\pm$187 1,425$\pm$73 $1.86\times 10^{6}\pm 4.51\times 10^{5}$ 82.25$\pm$9.30 6,278$\pm$2,245 924$\pm$251
GIRIL $2.15\times 10^{6}\pm 8.98\times 10^{5}$ 95.96$\pm$0.79 3,466$\pm$1,018 909$\pm$382 $5.20\times 10^{5}\pm 1.33\times 10^{5}$ 25.08$\pm$3.93 1,139$\pm$25 821$\pm$112 $4.33\times 10^{6}\pm 2.63\times 10^{5}$ 67.32$\pm$6.16 6,992$\pm$915 2,588$\pm$253
AIRL $3.11\times 10^{6}\pm 1.72\times 10^{5}$ 83.57$\pm$9.65 5,183$\pm$1,735 1,408$\pm$700 $2.53\times 10^{6}\pm 2.56\times 10^{5}$ 70.53$\pm$2.30 4,280$\pm$326 1,437$\pm$138 $2.31\times 10^{6}\pm 1.15\times 10^{5}$ 71.47$\pm$7.59 7,661$\pm$705 1,308$\pm$144
Max-Ent IRL $1.12\times 10^{6}\pm 3.54\times 10^{5}$ 85.48$\pm$10.65 2,570$\pm$710 523$\pm$152 $1.80\times 10^{6}\pm 5.17\times 10^{4}$ 68.81$\pm$0.75 3,756$\pm$318 1,046$\pm$29 $1.82\times 10^{6}\pm 3.09\times 10^{5}$ 83.04$\pm$6.07 4,658$\pm$1,489 882$\pm$192
@@ Appendix H Tukey HSD Test Results
Table 9: Tukey HSD test of the QD Score for the QD-IRL comparisons (w/o vs w/ EBC) in Table 8. $+$ indicates a positive effect of EBC. The significant pairwise comparisons ($p$<=0.05) are boldfaced.

### v4

# Diversifying Policy Behaviors with Extrinsic Behavioral Curiosity

@@ Abstract
Imitation learning (IL) has shown promise in various applications (e.g. robot locomotion) but is often limited to learning a single expert policy, constraining behavior diversity and robustness in unpredictable real-world scenarios. To address this, we introduce Quality Diversity Inverse Reinforcement Learning (QD-IRL), a novel framework that integrates quality-diversity optimization with IRL methods, enabling agents to learn diverse behaviors from limited demonstrations. This work introduces Extrinsic Behavioral Curiosity (EBC), which allows agents to receive additional curiosity rewards from an external critic based on how novel the behaviors are with respect to a large behavioral archive. To validate the effectiveness of EBC in exploring policies with diverse behaviors, we evaluate our method on multiple robot locomotion tasks. EBC improves the performance of QD-IRL instances with GAIL, VAIL, and DiffAIL across all included environments by up to 185%, 42%, and 150%, even surpassing expert performance by 20% in Humanoid. Furthermore, we demonstrate that EBC is applicable to Gradient-Arborescence-based Quality Diversity Reinforcement Learning (QD-RL) algorithms, where it substantially improves performance and provides a generic technique for learning behavioral-diverse policies. The source code of this work is provided at https://github.com/vanzll/EBC.
@@ 1 Introduction
  • Our framework can potentially enhance any IRL application that requires learning diverse policies, and can also significantly improve existing SOTA QD-RL algorithm, thus providing a generic framework for future research of learning behavioral-diverse policies.
@@ 2.1 Quality Diversity Reinforcement Learning
Distinct from traditional policy optimization which aims to find a single policy to maximize the cumulative reward, Quality Diversity Reinforcement Learning (QD-RL) aims to find a set of high-performing and diverse policies. Parametrizing policies based on $\theta\in\Theta\subset\mathbb{R}^{n}$, the objective function (i.e., fitness) is defined as $f:\Theta\rightarrow\mathbb{R}$, which is the cumulative reward in RL context. The diversity is based on a $k$-dimensional measure function $m:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$. The goal is to find policies $\theta\in\Theta$ for each local region in the behavior space $B=\mathbf{m}(\Theta)$. In grid-based algorithms MAP-Elites, one discretizes $B$ into $M$ cells (local regions) and forms a policy archive $\mathcal{A}$ , where each cell $i=1,\dots,M$ represents a small hypercube $\mathbf{a_{i}},\mathbf{b_{i}}]$ within a multi-dimensional grid of the behavior space. By filling new cells and replacing existing solutions if they are outperformed, the algorithm gradually evolves an archive of high-quality and diverse policies (Chatzilygeroudis et al., 2021; Pugh et al., 2016; Mouret & Clune, 2015a, b). Based on MAP-Elites, Co- variance Matrix Adaptation MAP-Elites (CMA-ME) (Fontaine et al., 2020) further adopts covariance matrix adaptation as the evolution algorithm to find new high-performing solutions more effectively and efficiently. In such algorithms, QD-RL can be treated equivalently to solving $M$ policy-optimization problems of the form
@@
$\text{s.t.}\quad\mathbf{m}(\theta)\in[\mathbf{a_{i}},\mathbf{b_{i}}],$
@@ 3 Extrinsic Behavioral Curiosity for QD-IRL
3:for $\text{iter}\leftarrow 1$ to NN do
4:  $f,g,\nabla g,\mathbf{m},\nabla\mathbf{m}\leftarrow\text{VPPO.jacobian}(\theta_{\mu},\mathcal{R},\mathbf{m},N_{1})$ $\triangleright$ approx grad using $\mathcal{R}$
@@
6:  _\_ $\leftarrow\text{update\_archive}(\theta_{\mu},f,\mathbf{m})$
@@
8:   $c\sim\mathcal{N}(\mu,\Sigma)$ $\triangleright$ sample gradient coefficients
@@
12:   $\Delta_{i}\leftarrow\text{update\_archive}(\theta_{i}^{\prime},f^{\prime},\mathbf{m}^{\prime})$ $\triangleright$ get archive improvement of each solution.
@@
16:  $g^{\prime}(\theta_{\mu}):=c_{\mu,0}g(\theta_{\mu})+\sum_{j=1}^{k}c_{\mu,j}m_{j}(\theta_{\mu})$, where $c_{\mu}=\mu^{\prime}$ $\triangleright$ construct EBC-QD objective function
17:  $\theta_{\mu}^{\prime}\leftarrow\text{VPPO.train}(\theta_{\mu},g^{\prime},N_{2},\mathcal{R})$ $\triangleright$ walk search policy using reward model $\mathcal{R}$
18:  $\mathcal{R}.\text{update}(\mathcal{D},\theta_{\mu}^{\prime})$ $\triangleright$ update reward model
@@ 3.3.3 DiffAIL-EBC
[…]
@@ 4.1 Learning Diverse High-Performing Skills from Limited Demonstrations
Experiments on Various Measure Metrics To further assess the generality of our framework, we applied EBC to the GAIL baseline (GAIL-EBC) and evaluated it on tasks that differ in both the type and the number of dimension of the QD measure $\mathbf{m}$. Specifically, we used: 1) Jump (dim=1dim{=}1), a scalar that records the height of the lowest foot at each step; 2) Angle (dim=2dim{=}2), a 2-D vector $\phi_{t}=(\cos\alpha,\ \sin\alpha)^{\top}$ capturing the yaw orientation $\alpha$ of the torso about the $z$-axis; and 3) Feet Contact (dim=4dim{=}4), a 4-D vector that stores, for each of the ant’s four legs, the fraction of time steps during which the leg touches the ground. Table 1 reports that GAIL-EBC improves QD-Score of GAIL throughout every scenario and significantly improve the Coverage of GAIL in humanoid environment.
@@
Environment Measure (dimdim) Model QD-Score Coverage (%)
@@
[…]
Ant Feet Contact (4) GAIL -484 468 $\pm$ 3 264 75.84 $\pm$ 0.0
[…]
@@ 5 Conclusion
In this work, we propose Extrinsic Behavioral Curiosity (EBC), a simple but effective approach which is highly extendable, and theoretically derived the rationale behind it. EBC can be seamlessly integrated into and can potentially improve any IRL method in QD contexts, providing the a generic QD-IRL framework for future research. Our framework follows the paradigm of IRL to learn a QD-enhanced reward function, and uses a QD-RL algorithm to optimize a policy archive. By encouraging behavior-level exploration, our framework addresses the key challenges of QD-IRL. Extensive experiments show that our framework can often improve QD-IRL methods and significantly outperform baselines. Additionally, we empirically prove that EBC can also significantly improve the performance of the state-of-the-art QD-RL algorithm, PPGA, highlighting the sub-optimality of the true reward function in QD-RL settings. With EBC, Gradient-Arborescence based QD-RL and QD-IRL algorithms can learn archives of diverse and high-performing policies more efficiently, pushing the boundary of learning behavioral-diverse policies. We also provide potential limitations of our work in Appendix K and additional discussion in Appendix L.
@@ E.2 Hyperparameters for IL
Variance schedule ($\beta_{t}$) linear from 1e41e-4 to 2e22e-2
@@ Appendix G QD Metrics Performance Results
[…]
@@ Appendix H Tukey HSD Test Results
Table 9: Tukey HSD test of the QD Score for the QD-IRL comparisons (w/o vs w/ EBC) in Table 8. ++ indicates a positive effect of EBC. The significant pairwise comparisons (pp<=0.05) are boldfaced.